A man invested $100,000. Part of it is invested at 10% and the rest at 8%. If the total income is $8800, how mush did he invest at 10%?
Answer by mananth(16946)

Investment I 10.00% per annum ----x
Investment II 8.00% per annum ----y

x+y= 100000 ------------------------1
10.00% x+8.00% y=$8,800.00
Multiply by 100
10x+8y= $880,000.00 --------2

Multiply (1) by -10
we get

-10x-10 y =-1000000.00

Add this to (2)

-2 y = -$120,000.00

divide by -2

y = $60,000.00 investment at 8.00%
Balance $40,000.00 investment at 10.00%
CHECK
$40,000.00 $4,000.00
$60,000.00 $4,800.00
Total $8,800.00